3. Your firm’s accountant believes that 15% of the company’s invoices contain arithmetic errors. To check this theory, the accountant randomly selects 50 and checks them.
 
a) Give the formula for the probability that exactly six of them are in error.
b) Give the mean and standard deviation for distribution representing the number of invoices among the 50 that are in error.
c) Which is larger: the mean or the median? Why?

Answer by stanbon(75887)

Your firm’s accountant believes that 15% of the company’s invoices contain arithmetic errors. To check this theory, the accountant randomly selects 50 and checks them.

a) Give the formula for the probability that exactly six of them are in error
P(x = 6) = 50C6*(0.15)^6*(0.85)^45
------------------------------------------------
b) Give the mean and standard deviation for distribution representing the number of invoices among the 50 that are in error.
mean = np = 50*0.15 = 7.5
std = sqrt(npq) = sqrt(7.5*9.85) = 2.5249
------
c) Which is larger: the mean or the median? Why?
median = middle score of 0 to 15::: (7+8)/2 = 7.5
mean = np = 7.5
----
The mean = the median
=========================
Cheers,
Stan H.